# On $3$-manifolds admitting co-orientable taut foliations, but none with vanishing Euler class

### v2

@@ Abstract.
Key words: taut foliation, Knot Floer complex, self-conjugate spinc structure, Seifert fibred manifold, normal bundle, Euler class, horizontal foliation.
@@ 1. Introduction
1.1. Rational homology spheres admitting no taut foliations with vanishing Euler class
It is predicted by the $L$-space conjecture [4, 18] that if a closed, orientable, prime $3$-manifold $M$ admits a co-orientable taut foliation, then its fundamental group is left-orderable. One of the main tools for proving this is to use Thurston’s universal circle action, which to each co-oriented taut foliation $\mathcal{F}$ on an oriented rational homology $3$-sphere $M$ associates a non-trivial representation
@@ 1.1. Rational homology spheres admitting no taut foliations with vanishing Euler class
See [10] for the details. If one can show that this representation lifts to the universal covering group $\mbox{Homeo}_{\mathbb{Z}}(\mathbb{R})$ of $\mbox{Homeo}_{+}(S^{1})$, then it follows from [8, Theorem 1.1] that the fundamental group of $M$ is left-orderable. However, such a lift exists if and only if the Euler class of $\rho_{\mathcal{F}}$ vanishes. Since this Euler class coincides with the Euler class of the tangent bundle of $\mathcal{F}$ [5], which we denote by $e(\mathcal{F})$, we are naturally led to the following question.
Question 1.1.
If a closed, orientable, prime $3$-manifold admits a co-orientable taut foliation, does it admit one whose Euler class is zero?  [^1] Given a co-orientable taut foliation, the Euler class of its tangent bundle is well-defined up to sign. So it makes sense to discuss if a co-orientable taut foliation has vanishing Euler class.
Although it has been expected that the answer to the question is no, we do not know of any examples verifying this. In this article, we show that such examples occur quite frequently.
Convention: When referring to $p/q$-surgery on a knot, we assume that $p$ and $q$ are coprime integers with $p\geq 0$.
@@
Let $M$ be obtained by $p/q$-surgery on a knot $K$ in the $3$-sphere. If $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1, then $M$ does not admit a co-orientable taut foliation with Euler class zero.
Thus the existence of a knot for which some large surgery with even denominator admits a co-orientable taut foliation shows that the answer to Question 1.1 is no. For instance, Roberts has shown that if $K$ is a non-trivial fibred knot in the $3$-sphere, then $K(p/q)$, the manifold obtained by $p/q$-surgery on $K$, admits a co-orientable taut foliation for all $p/q$, $p/q<1$, or p/q>1p/q>-1 depending on whether the fractional Dehn twist coefficient of $K$ is zero, positive, or negative respectively [26]. The following corollary states the case when the fractional Dehn twist coefficient is zero. Examples of such knots include all hyperbolic fibred alternating knots [6, 2].
@@
Suppose that $K$ is a non-trivial fibred knot in the $3$-sphere whose fractional Dehn twist coefficient is zero. Then for all slopes $p/q$ with $q$ even and |p/q|>2g(K)1|p/q|>2g(K)-1, $K(p/q)$ admits a co-orientable taut foliation, but none with zero Euler class.
A similar conclusion holds for many non-fibred knots. Indeed, the $L$-space conjecture predicts that every irreducible rational surgery on a non-fibred knot admits a co-orientable taut foliation and this is known to be true in many instances. For instance, Roberts and Delman proved this for non-torus alternating knots and for Montesinos knots which are not $L$-space knots.
Examples of manifolds having left-orderable fundamental groups and admitting co-orientable taut foliations, but none with vanishing Euler class, can also be constructed, thus showing that no left-orders on the group arise from lifts of universal circle representations. For instance, the figure eight knot is a fibred hyperbolic knot of genus one with zero fractional Dehn twist coefficient and each rational surgery on it has a left-orderable fundamental group [27].
@@
Suppose that $K$ is the figure eight knot. Then for all slopes $p/q$ with $q$ even and $|p/q|>1$, $K(p/q)$ has a left-orderable fundamental group and admits a co-orientable taut foliation, but no co-orientable taut foliation with zero Euler class.
By way of contrast, we note that Theorem 1.2 cannot be extended to all large surgeries on a knot without some hypotheses on the slope $p/q$. For instance, it is known that $p$-surgery on the figure eight knot admits co-orientable taut foliations of zero Euler class for each integer $p$ [11, 14]. More generally, we have
@@
Let $M$ be a rational homology sphere obtained by integer surgery on a genus one knot in the $3$-sphere. Then every co-orientable taut foliation on $M$ has Euler class zero.
Remark 1.5.
The line of reasoning used in the proof of Proposition 2.3 can easily be extended to show that if $K$ is an almost $L$-space knot, meaning a knot with positive almost $L$-space surgeries [9], and $M$ is obtained by $p$-surgery on $K$ for some $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero. This shows that for knots of arbitrarily large genus, constraints must be placed on the denominators of surgery slopes for Theorem 1.2 to hold.
Corollary 1.6.
Suppose that $K$ is a knot of genus one in the $3$-sphere. If $p$ is an integer such that $K(p)$ admits a co-orientable taut foliation, then $\pi_{1}(K(p))$ is left-orderable. Similarly, if $K$ is an almost $L$-space knot and $p\geq 2g(K)-1$ is an integer such that $K(p)$ admits a co-orientable taut foliation, then $\pi_{1}(K(p))$ is left-orderable.
All the examples in Corollary 1.4 are hyperbolic. The simplest examples of Seifert fibred manifolds in which every co-orientable taut foliation has non-zero Euler class arise from Dehn surgery on the trefoil knot.
Theorem 1.7.
Suppose that $K$ is the right-handed trefoil knot. Then $K(p/q)$ admits a co-orientable taut foliation and has a left-orderable fundamental group if and only if $p/q<1$. It admits a co-orientable taut foliation with zero Euler class if and only if $p/q<1$ and $|q|\equiv 1\pmod{p}$.
In general, we denote a closed oriented Seifert fibred manifold $M$ with orientable base orbifold by $M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, where $b\in\mathbb{Z}$ and $b_{i}\geq 2$ for $i=1,\ldots,n$. See §3.1 for a detailed explanation of the notation. We denote the base orbifold by $\mathcal{B}$; this is a closed orientable surface $|\mathcal{B}|$ of genus $g\geq 0$ with $n\geq 0$ cone points of orders $b_{1},...,b_{n}$. Recall that the Euler characteristic $\chi(\mathcal{B})$ of $\mathcal{B}$ is $\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$, and the Euler number of $M$ is $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$.
@@
We can also produce toroidal examples which are not Seifert fibred. Define the JSJ graph of a knot $K$ to be the rooted tree dual to the JSJ tori of the exterior $X(K)$ of $K$, where the root vertex corresponds to the piece of the JSJ decomposition containing $\partial X(K)$. Any finite rooted tree can arise as the JSJ graph of a knot. We say that a JSJ graph is a rooted interval when it is homeomorphic to an interval with root at an endpoint. Since non-integral surgeries on knots in $S^{3}$ are irreducible [12], the following corollary follows immediately from Theorem 1.2 and Theorems 3.13 and 4.3 of [3].
@@
Remark 1.10.
1.2. Euler class of the normal bundle of a Seifert fibration
While Theorem 1.2 is proved using the knot Floer complex, the key fact in analyzing the Euler classes of co-oriented taut foliations on a Seifert fibred rational homology $3$-sphere $M$ is that, up to isotopy, all such foliations are horizontal; that is, transverse to the Seifert fibres of $M$ (Lemma 5.1). Hence, the tangent bundle of $\mathcal{F}$ is isomorphic to the normal bundle $\nu_{M}$ to the Seifert fibres of $M$, while the latter makes sense even without the existence of a horizontal foliation.
We will refer to $\nu_{M}$ as the normal bundle of the Seifert fibration. In §3 and §4, we determine the following necessary and sufficient conditions for its Euler class to vanish.
Theorem 1.11.
@@ 1.2. Euler class of the normal bundle of a Seifert fibration
    $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n,\;\text{and}$
  2. (2)
    $me(M)=\chi(\mathcal{B})$.
Theorem 1.8 follows from Theorem 1.11 and Lemma 5.1.
Next, we deduce the following easily verifiable necessary condition for $\nu_{M}$ to be a trivial bundle:
Corollary 1.12.
Given a closed oriented Seifert fibred manifold $M$, if the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes, then either $M$ is an orientable $S^{1}$-bundle over the torus or one of the following holds:
@@
Note that if $M$ is an orientable $S^{1}$-bundle over the torus, then its normal bundle is always trivial (see Remark 3.2).
By abelianizing the presentation (4.0.1) of $\pi_{1}(M)$ obtained by viewing $M$ as doing a Dehn filling of a trivial $S^{1}$-bundle over an $n$-punctured orientable surface of genus $g$ as in §3.1, one can show that if $T_{1}(M)$ denotes the torsion subgroup of $H_{1}(M)$, then
@@
where we take the convention that $b_{1}b_{2}\cdots b_{n}=1$ when $n=0$. Thus, it follows from Corollary 1.12 that:
@@
Consider the Seifert fibred manifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ and suppose that the Euler class of the normal bundle $\nu_{M}$ of the Seifert fibration vanishes. Then the integer $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is a multiple of $|T_{1}(M)|$. Hence, if $M$ is a rational homology $3$-sphere, then $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is a multiple of $|H_{1}(M)|$.
We will present more examples of rational homology Seifert fibred manifolds that admit co-orientable taut foliations, but none with vanishing Euler class, in §5.3 (Examples 5.3). In fact, our computations in §5 suggest that most closed Seifert fibred manifolds have nontrivial normal bundle, including the case when the base orbifold is Euclidean (i.e. $\chi(\mathcal{B})=0$). This may seem counterintuitive, since we have pointed out that the normal bundle is always trivial when the base $\mathcal{B}$ is the torus.
We close with an interesting by-product of our analysis.
@@
In contrast, closed Seifert fibred manifolds with spherical base orbifolds have finite fundamental groups and hence cannot admit taut foliations. Yet, in §5.1, we will show examples of such manifolds for which $e(\nu_{M})$ vanishes.
@@
We prove Theorem 1.2 and Proposition 2.3 in §2. In §3 we explain our notation for Seifert manifolds and prove Theorem 1.11, which gives necessary and sufficient conditions for the Euler class of the normal bundle of a Seifert fibration to vanish. We also prove Corollary 1.12 in this section. In §4 we present a different perspective on Theorem 1.11 and Corollary 1.12 in the case that the base orbifold is hyperbolic through the use of discrete faithful $PSL(2,\mathbb{R})$-representations of $\pi_{1}(\mathcal{B})$. As a consequence of this approach, we prove the hyperbolic case of Theorem 1.14. Finally, in §5 we prove Theorems 1.7 and 1.8, and provide examples of Seifert fibred manifolds whose normal bundles are (non-)trivial. We classify which occurs when $\mathcal{B}$ is spherical in §5.1 and is Euclidean in §5.2. The Euclidean case of Theorem 1.14 is dealt with at the end of §5.2. The case that $\mathcal{B}$ is hyperbolic is analysed in §5.3.

### v3

@@ Abstract.
Key words: taut foliation, Euler class, knot Floer complex, self-conjugate spinc structure, Seifert fibred manifold, normal bundle, horizontal foliation.
@@ 1. Introduction
A long-standing problem in $3$-manifold topology is to determine which classes in the second cohomology group of a closed, oriented $3$-manifold $M$ are the Euler classes of tangent bundles of co-oriented taut foliations on $M$. Such classes are constrained by the fact that every orientable $3$-manifold is parallelizable, which implies that the Euler class of a co-oriented plane field over $M$ is an even class (see [18, Proposition 2.3] for instance). If we drop the tautness condition, every even class is realised, since each co-oriented plane field on $M$ is homotopic to an integrable one [43]. This no longer holds when one requires that the foliation be taut. Indeed, only finitely many homotopy classes of co-oriented plane fields can be realised as the tangent plane field of a taut foliation [25, Corollary 1.5], [12, Corollary 6.18]. Determining which cohomology classes can occur as Euler classes of co-orientable taut foliations remains open.
Much work on the problem has been done when $b_{1}(M)>0$. In this case, Gabai has shown that taut foliations always exist [11]. Further, results of Roussarie and Thurston show that the dual Thurston norm of the Euler class of any co-oriented taut foliation is at most one [41, 44]. Thurston also conjectured that any class in $H^{2}(M)$ of dual Thurston norm one can be realised as the Euler class of a co-oriented taut foliation. However, this conjecture is now known to be false, following counterexamples constructed by Yazdi and Liu [47, 28]. At the other extreme, one may ask whether the zero class can be realised as the Euler class of a co-oriented taut foliation. See [47, Question 9.1 and Question 9.2] and the related discussions therein. Also see [10].
In this article, we consider the case $b_{1}(M)=0$; that is, $M$ is a rational homology sphere or equivalently, $H^{2}(M)$ is finite. This setting differs from the case $b_{1}(M)>0$ in two fundamental ways. First, the existence of co-orientable taut foliations is no longer guaranteed. Second, the dual Thurston norm vanishes, since $H^{2}(M)$ is finite, so the usual techniques used to analyze the Euler class when $b_{1}(M)>0$ do not apply. Nevertheless, the zero class remains a distinguished element of $H^{2}(M)$. This leads naturally to the following question (cf. [47, Question 9.2]):
Question 1.1.
If a closed, orientable rational homology sphere admits a co-orientable taut foliation, does it admit one whose Euler class is zero?  [^1] Given a co-orientable taut foliation, the Euler class of its tangent bundle is well-defined up to sign. So it makes sense to discuss if a co-orientable taut foliation has vanishing Euler class.
Question 1.1 takes on further significance in the context of the $L$-space conjecture [3, 24], which predicts that if a closed, orientable, prime $3$-manifold $M$ admits a co-orientable taut foliation, then its fundamental group is left-orderable. The conjecture is known to hold when $b_{1}(M)>0$ [11, 7]. On the other hand, when $M$ is a rational homology sphere, one of the main tools for proving this is to use Thurston’s universal circle action, which to each co-oriented taut foliation $\mathcal{F}$ on an oriented rational homology $3$-sphere $M$ associates a non-trivial representation
@@
See [9] for the details. If one can show that this representation lifts to the universal covering group $\mbox{Homeo}_{\mathbb{Z}}(\mathbb{R})$ of $\mbox{Homeo}_{+}(S^{1})$, then it follows from [7, Theorem 1.1] that the fundamental group of $M$ is left-orderable. However, such a lift exists if and only if the Euler class of $\rho_{\mathcal{F}}$ vanishes. Since this Euler class coincides with the Euler class of the tangent bundle of $\mathcal{F}$ [5], which we denote by $e(\mathcal{F})$, we are naturally led to Question 1.1
For the reasons discussed above, Question 1.1 has been raised often over the years. Its answer is yes when $H^{2}(M)$ is a (possibly trivial) group consisting of 2-torsion, since the only even class is zero in this case. And though it would be surprising if the answer were yes in general, no examples of manifolds for which the answer is no have appeared in the literature. In this article, we show that such examples are abundant, and ubiquitous amongst manifolds obtained by surgery on knots in the $3$-sphere (Theorem 1.7) and amongst Seifert fibred rational homology $3$-spheres (Theorem 1.13). Further, many of our examples have left-orderable fundamental groups, even though no left-order on the group arises from a lift of a universal circle representation. Thus we ask the following question.
Question 1.2.
Is there a general procedure for constructing, from a closed, orientable, prime 3-manifold which admits a co-orientable taut foliation, a non-trivial action of its fundamental group on the reals by orientation-preserving homeomorphisms?
1.1. Surgery on knots in the $3$-sphere
The examples derived from surgery are obtained by applying recent work of Lin, who showed that if a $3$-manifold $M$ admits a co-oriented taut foliation $\mathcal{F}$, then there is a $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ on $M$ whose associated plane field is $T\mathcal{F}$ and a direct $\mathbb{F}$ summand of the $\mathbb{F}[U]$-module $HF_{\mathrm{red}}(M,\mathfrak{s})$ [27].
Convention: When referring to $p/q$-surgery on a knot, we assume that $p$ and $q$ are coprime integers with $p>0$. We denote the manifold obtained by $p/q$-surgery on a knot $K$ by $K(p/q)$.
@@ 1.1. Surgery on knots in the $3$-sphere
Let $K$ be a knot in the $3$-sphere. If $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1, then $K(p/q)$ does not admit a co-orientable taut foliation with Euler class zero.
In [18, Theorem 1.12(1)], it was shown that if $|p/q|>2g(K)$ and $p/q$ is not an integer, then any co-orientable taut foliation on $K(p/q)$ that is transverse to the core of the surgery solid torus cannot have vanishing Euler class. The above theorem removes the transversality condition under the assumption that $q$ is even.
Thus by Theorem 1.3, the existence of a knot for which some large surgery with even denominator admits a co-orientable taut foliation shows that the answer to Question 1.1 is no. We give some examples in the corollary below.
@@
Suppose that $K$ is a non-trivial knot in the $3$-sphere which is either
  1. (1)
    fibred with zero fractional Dehn twist coefficient, or
  2. (2)
    alternating but not a $(2,q)$-torus knot, or
  3. (3)
    a Montesinos knot but not a $(2,q)$-torus knot or a (2,3,2n+1)(-2,3,2n+1)-pretzel knot.
Then for all slopes $p/q$ with $q$ even and |p/q|>2g(K)1|p/q|>2g(K)-1, $K(p/q)$ admits a co-orientable taut foliation, but none with zero Euler class.
Part (1) of Corollary 1.4 follows from Theorem 1.3 together with Roberts’ result [40], while parts (2) and (3) follow from Theorem 1.3 and work of Delman and Roberts.
As we mentioned above, examples of manifolds having left-orderable fundamental groups and admitting co-orientable taut foliations, but none with vanishing Euler class, can also be constructed, thus showing that no left-orders on the group arise from lifts of universal circle representations. Here are the simplest examples. More examples are provided in Theorem 1.7 and Theorem 1.13 below.
  • The figure eight knot $K$ is a fibred hyperbolic knot of genus one with zero fractional Dehn twist coefficient and each rational surgery on it has a left-orderable fundamental group [48]. Hence, Corollary 1.4 implies that for all slopes $p/q$ with $q$ even and $|p/q|>1$, $K(p/q)$ has a left-orderable fundamental group and admits a co-orientable taut foliation, but no co-orientable taut foliation with zero Euler class.
  • Suppose that $K$ is the right-handed trefoil knot. Then $K(p/q)$ admits a co-orientable taut foliation (equivalently, has a left-orderable fundamental group [7]) if and only if $p/q<1$ [23, 32]. It follows from Lemma 3.5 and [18, Theorem 1.7] that $K(p/q)$ admits a co-orientable taut foliation with zero Euler class if and only if $p/q<1$ and $|q|\equiv 1\pmod{p}$.
All the examples in the first family are hyperbolic, while those in the second are Seifert fibred. We can also use Theorem 1.3 to produce toroidal examples which are not Seifert fibred.
Define the JSJ graph of a knot $K$ to be the rooted tree dual to the JSJ tori of the exterior $X(K)$ of $K$, where the root vertex corresponds to the piece of the JSJ decomposition containing $\partial X(K)$. Any finite rooted tree can arise as the JSJ graph of a knot. We say that a JSJ graph is a rooted interval when it is homeomorphic to an interval with root at an endpoint. Since non-integral surgeries on knots in $S^{3}$ are irreducible [14], the following corollary follows immediately from Theorem 1.3 and Theorems 3.13 and 4.3 of [2].
@@
Remark 1.6.
If $K$ is an $L$-space knot then for slopes $p/q\geq 2g(K)-1$ the manifold $K(p/q)$ is an $L$-space and hence does not admit a co-orientable taut foliation. On the other hand, since $L$-space knots are fibred, Roberts’ work [40] shows that $K(p/q)$ has a co-orientable taut foliation if $p/q<1$. We establish conditions under which negative surgeries on an $L$-space knot cannot admit a co-orientable taut foliation with Euler class zero. These conditions are phrased in terms of congruence relations involving the torsion coefficients of the Alexander polynomial (Theorem 2.8 and Corollary 2.9), as well as the bridge index (Corollary 2.11).
By performing negative surgeries on the (2,3,2n+1)(-2,3,2n+1)-pretzel knots, Theorem 2.8 provides the following examples.
Theorem 1.7.
For any integer $p\geq 7$ there are infinitely many hyperbolic 3-manifolds $M$ such that $H^{2}(M)\cong\mathbb{Z}/p$ and $M$ admits a co-orientable taut foliation, but none with Euler class zero. Further, $\pi_{1}(M)$ is left-orderable.
By way of contrast, note that Theorem 1.3 cannot be extended to large surgeries on all knots without some hypotheses on the slope $p/q$. For instance, if $M$ is a rational homology sphere for which there is a unique $\mathrm{spin}^{c}$-structure $\mathfrak{s}_{0}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})$ is non-zero, then any co-orientable taut foliation on $M$ has Euler class zero (see Proposition 2.4). As a consequence, we deduce,
@@
Let $M$ be a rational homology sphere obtained by $p$-surgery on a knot $K$ is the $3$-sphere, for some $p\in\mathbb{Z}$. If $K$ has genus one, or $K$ is an almost $L$-space knot  [^2] KK is an almost $L$-space knot if it has a positive almost $L$-space surgery of slope $p/q\geq 2g(K)-1$ [8]. and $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero.
Finally, recalling the theorem of Hu from [18] mentioned above (also see Corollary 2.9 and Remark 2.10), we ask the following question.
Question 1.8.
In Theorem 1.3, can the assumption that $q$ be even be replaced by the assumption $|q|>1$?
1.2. Seifert fibred manifolds
We denote a closed oriented Seifert fibred manifold $M$ with orientable base orbifold by $M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, where $b\in\mathbb{Z}$ and $b_{i}\geq 2$ for $i=1,\ldots,n$. See §3.1 for a detailed explanation of the notation. We denote the base orbifold by $\mathcal{B}$; this is a closed orientable surface $|\mathcal{B}|$ of genus $g\geq 0$ with $n\geq 0$ cone points of orders $b_{1},...,b_{n}$. Recall that the Euler characteristic $\chi(\mathcal{B})$ of $\mathcal{B}$ is $\chi(|\mathcal{B}|)-n+\sum_{i=1}^{n}\frac{1}{b_{i}}$, and the Euler number of $M$ is $e(M)=b+\sum_{i=1}^{n}\frac{a_{i}}{b_{i}}$.
A Seifert fibred rational homology $3$-sphere $M$ which admits a co-orientable taut foliation $\mathcal{F}$ is necessarily of the form $M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ [1, Proposition 6.1]. Further, $\mathcal{F}$ is necessarily horizontal, that is, transverse to the Seifert fibres of $M$ (see Lemma 3.5). As such, the tangent bundle of $\mathcal{F}$ is isomorphic to the normal bundle $\nu_{M}$ to the Seifert fibres of $M$. In §3 we determine necessary and sufficient conditions for the Euler class of $\nu_{M}$ to vanish (Theorem 3.3), which leads to the following theorem.
@@ 1.2. Seifert fibred manifolds
Given a closed oriented Seifert fibred manifold $M$ which admits a co-orientable taut foliation whose Euler class vanishes, then one of the following holds:
@@
    $M$ is an orientable $S^{1}$-bundle over the torus,
@@
By abelianizing the presentation (4.0.1) of $\pi_{1}(M)$ obtained by viewing $M$ as doing a Dehn filling of a trivial $S^{1}$-bundle over an $n$-punctured orientable surface of genus $g$ as in §3.1, one can show that if $e(M)\neq 0$ and $T_{1}(M)$ denotes the torsion subgroup of $H_{1}(M)$, then
@@
where we take the convention that $b_{1}b_{2}\cdots b_{n}=1$ when $n=0$. Thus, it follows from Theorem 1.9 that:
@@
Suppose that $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ is a rational homology $3$-sphere which admits a co-orientable taut foliation $\mathcal{F}$ with zero Euler class. If $e(M)\neq 0$ then the integer $b_{1}\ldots b_{n}\chi(\mathcal{B})$ is an integer multiple of $|H_{1}(M)|$.
When the base orbifold $M=M\bigl(g;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$ is hyperbolic, we provide a second proof of Theorem 1.9 from the viewpoint of discrete faithful representations of Fuchsian groups. Here is an interesting by-product of the analysis.
@@
As an application of Theorem 1.9 we show that such examples are actually abundant in the presence of higher order torsion.
Theorem 1.13.
Let $G$ be a finite abelian group containing an element of order at least three. Then there exist infinitely many Seifert fibred manifolds $M$ such that $H_{1}(M)\cong G$ and $M$ admits a co-orientable taut foliation, but none with Euler class zero. Further, $\pi_{1}(M)$ is left orderable.
Remark 1.14.
It is known that Seifert fibred manifolds satisfy the $L$-space conjecture [7, 30, 3], so all such manifolds which admit co-orientable taut foliations have left-orderable fundamental groups even though it is often the case that this cannot be shown using universal circle representations. One uses, rather, the fact that co-orientable taut foliations on Seifert fibred rational homology spheres are $\mathbb{R}$-covered.
We will present more examples of Seifert fibred manifolds that admit co-orientable taut foliations, but none with vanishing Euler class, in §5.3. In fact, our computations in §5 suggest that most closed Seifert fibred manifolds have nontrivial normal bundle, including the case when the base orbifold is Euclidean (i.e. $\chi(\mathcal{B})=0$). This may seem counterintuitive, since we have pointed out that the normal bundle is always trivial when the base $\mathcal{B}$ is the torus. In contrast, closed Seifert fibred rational homology spheres with spherical base orbifolds have finite fundamental groups and hence cannot admit taut foliations. Yet, in §5.1, we will construct examples of such manifolds for which $e(\nu_{M})$ vanishes.
@@
In §2, we apply Heegaard Floer homology to produce hyperbolic $3$-manifolds obtained by Dehn filling on knots in the $3$-sphere that admit co-orientable taut foliations, but none with vanishing Euler class. All results stated in §1.1 are proved in this section. In §3 we explain our notation for Seifert manifolds and prove Theorem 3.3, which characterizes those Seifert manifolds for which normal bundle $\nu_{M}$ to the Seifert fibres is trivial. Theorem 1.9 then follows from this and Lemma 3.5. In §4 we study the Euler class of the normal bundle $\nu_{M}$ through the use of discrete faithful $PSL(2,\mathbb{R})$-representations of $\pi_{1}(\mathcal{B})$ in the case that the base orbifold $\mathcal{B}$ is hyperbolic. As a consequence of this approach, we provide a different proof of Theorem 3.3 when the base is hyperbolic and prove the hyperbolic case of Theorem 1.12. Finally, in §5 we provide examples of Seifert fibred manifolds whose normal bundles are (non-)trivial. We classify which occurs when $\mathcal{B}$ is spherical in §5.1 and is Euclidean in §5.2. The Euclidean case of Theorem 1.12 is dealt with at the end of §5.2. The case that $\mathcal{B}$ is hyperbolic is analysed in §5.3.